Referee of the submission

“COMPLEX SYSTEMS DYNAMICS IN AN ECONOMIC

MODEL WITH MEAN FIELD INTERACTIONS ” by Gianfranco Giulioni

This paper deals the simulated dynamics of a simple agent based economic system are analyzed. These dynamics are of the complex systems type in the sense that the degree of selforganization changes with time. The paper is well written and the link with Complex Systems clearly outlined. I strongly suggest to accept the contribution.